Ophel has Php53.00 in 5 peso and 1 peso coins pieces. If she has 25 coins all, how many has she of each kind?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Ophel has Php53.00 in 5 peso and 1 peso coins pieces. If she has 25 coins all, how many has she of each kind?
----------------
How many pesos in a Php?

Answer by ikleyn(52788)

.

     x +  y = 25      (1)    (counting coins)

    5x + 1y = 53      (2)    (counting money)



Subtract equation (1) from equation (2) and solve for x


     4x     = 53 - 25 = 28

      x               = 28/4 = 7.


ANSWER.  7 5-peso coins and the rest, 25-7 = 18 1-peso coins.


CHECK.   7*5 + 18*1 = 35 + 18 = 53 Php.    ! Precisely correct !

Solved.